OpenStudy (gazela1997):

factor 0=x^2=7x-30 MEDAL AND FAN

OpenStudy (gazela1997):

OOO ITS 0=x^2+7x-30

OpenStudy (anonymous):

when you move over the 7x + 30 would their signs not change? so wouldnt it be 0= x^2 - 7x +30?

OpenStudy (muzzack):

Solution Steps 1. Swap Both Sides of Equation x^2+7*x−30=0 2. Multiply Terms x^2+7x−30=0 3. Move Constants to Right x^2+7x=0+30 4. Remove Zero Terms x^2+7x=30 5. Factor Trinomial (x+10)*(x−3)=0 Forked Equation 1 x+10=0 Forked Equation 2 x−3=0 Now Solving Equation 1 x+10=0 6. Move Constants to Right x=0−10 7. Remove Zero Terms x=-10 End of Equation 1 Now Solving Equation 2 x−3=0 8. Move Constants to Right x=0+3 9. Remove Zero Terms x=3 End of Equation 2 Plug candidate answers into original equation... Answer 1: -10 Answer 2: 3

OpenStudy (gazela1997):

holy pellet you just made more sense than my teacher did for 2 weeks

OpenStudy (muzzack):

lol thx ;) its easy